OpenStudy (anonymous):

*MEDAL* Graph this: -5x + 2y = 2

OpenStudy (anonymous):

Put it in the form y = mx + c where m would be a number times x (the gradient) and c would be point on which the graph intersects the y axis.

OpenStudy (anonymous):

Here's my choices:

OpenStudy (anonymous):

I'll tell you which one is right if you have a go!

OpenStudy (anonymous):

It would have to be the 2nd or 4th one,r ight?

OpenStudy (anonymous):

No, i have time :)

OpenStudy (anonymous):

It couldn't be the second one, try rearranging it, you'll see it straight away

OpenStudy (anonymous):

Hmm...

OpenStudy (anonymous):

Is it the 3rd one?

OpenStudy (anonymous):

y=(2+5x)/2 right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

So that simplifies to?

OpenStudy (anonymous):

3.5x?

OpenStudy (anonymous):

y = (5/2) x + 1

OpenStudy (anonymous):

2.5x + 1...?

OpenStudy (anonymous):

Yes, so the graph has to cross the y axis at 1, which one of your diagrams crosses the y axis at 1?

OpenStudy (anonymous):

Its the first one I think...

OpenStudy (anonymous):

The 1st one, yes.

OpenStudy (anonymous):

Yeah, it's the first one :)

OpenStudy (anonymous):

Thanks, dmot! :D

OpenStudy (anonymous):

No problem :)